Find the four fourth roots ofz = 5 + 11j
Answer #1

3= 5+ lli V4 We have to find (5+ c) W= reio (أ) Let w 4 = 5 tlli whole then (orelojes 5tlla gry i (48) = 5+ first we have to

If you have any doubt feel free to ask in the comment section.